You want to buy a car. The loan amount will be $24,000. The company is offering a 2% interest rate for 36 months (3 years). What will your monthly payments be?
1
Expert's answer
2020-05-08T19:45:53-0400

The monthly loan repayment will be computed using the following formula:

"\text{p}=\dfrac{D (1- (1+\dfrac{r}{k} )^{-nk} }{ \dfrac{r}{k}}"


Where p is the loan balance at the beginning of the period which is $24000.

d is the loan repayment.

r is the rate of interest which is 2%.

k is the number of compounding period in a year (12 months)

n is the length of the loan (3years)


"24000=\dfrac{D (1- (1+\dfrac{0.02}{12} )^{-12 \times3} }{ \dfrac{0.02}{12}}"


"24000=34.91306D"


"D=\dfrac{24000}{34.91306}"


"D= $687.42"

The monthly payment is $687.42